# Cross Validation in Stochastic Analytic Continuation

### v2

@@ I Introduction
Dynamic response functions (spectral functions) provide a key link between the quantum Monte Carlo (QMC) simulations of many-body systems and their experimental counterparts. While it is intractable to measure these spectral function directly, one can instead preform numerical analytic continuation of imaginary-time correlation functions to real frequency. However, this process is what is considered an “ill-posed” problem [1, 2, 3], meaning that there may be many viable solutions for the spectral function fitting a set of QMC-generated data. The Maximum Entropy method (MEM) [4, 5, 6, 7] has been the traditional tool used to tackle this problem and has produced many important results [8]. However, MEM has its limitations, being able only to reproduce broad spectral features, unless prior information about more complicated structure is provided to the algorithm.
@@
Cross validation has a natural extension to SAC. The imaginary-time correlation function data, calculated using QMC simulations, are typically stored as individual bins, which can be split up into sampling and validation sets. Here, the goodness-of-fit $\chi^{2}$ is used as the loss function to “train” and to “validate” the resulting spectral function. The concept of using cross validation within SAC was first explored by Efremkin et al. [24], whose work we draw on and significantly expand upon. Here, we preform an exhaustive study using both synthetic and real QMC-generated data.
In Sec. II we describe the SAC procedure generally, and in Sec. III we explain our cross validation scheme in detail. In Sec. IV.1 and  IV.2 we implement cross validation using the unconstrained SAC method and demonstrate how it can be used to determine the optimal sampling temperature, which we compare to the simple criterion proposed previously [20]. In Sec. V.1 and  V.2, we show how cross validation can be used in practice to determine which unconstrained SAC sampling parameterization produces the most statistically likely spectrum. We extend this model selection procedure to the class of constrained SAC parameterization in Sec. VI. In Sec. VI.1 we preform a test using QMC-generated data for the $S=1/2$ antiferromagnetic (AFM) Heisenberg model, where it is known that the spectral function for the operator $O=S_{q}^{z}$ contains a power-law divergent edge. In Sec. VI.2 we preform a preliminary test on the $S=1/2$ AFM Heisenberg model with long-range interactions [25, 26], a model where the exact features of the spectral function are unknown. We summarize our results and discuss the remaining open questions and future directions in Sec. VII.
@@ II SAC Procedure
$\bar{G}_{k}(\tau_{i})=\frac{1}{N_{k}}\sum_{b\in k}G^{b}(\tau_{i}).$ (6)
We define the “error level” for these sets of $G(\tau)$ data as the magnitude of the error bar on the averaged correlation function, $\bar{G}_{k}(\tau_{i})$, for the largest $\tau$ point included, when normalized so that $\bar{G}_{k}(0)=G^{b}(0)=1$. For larger $\beta$, a cutoff is typically introduced at the $\tau$ value where the relative error on $\bar{G}_{k}(\tau)$ is greater than 10%, which we would instead use to quantify the error level.
@@ IV.1 Single Gaussian Peak
in some cases with $A_{i}$ variable and sampled [11], as depicted in Fig. 1(b), but in this case we use fixed amplitudes $A_{i}=1/N_{\omega}$, Fig. 1(a). In this example, we use $N_{\omega}=5,000$ total $\delta$-functions, which is large enough to no longer detect and dependance of $S(\omega)$ on $N_{\omega}$.
@@ IV.2 Two Gaussian Peaks
We preformed the same test as in Sec. IV.1, and present the results in Fig. 4. In this case, both the validation minimum and the temperature range corresponding to the criterion with $a$ ranging from 0.25 and 1 are much broader, but still align quite well with one another. The higher level of uncertainty is a reflection of the inability of the unconstrained sampling scheme to reproduce this spectrum nearly as well as in the previous case, which can be seen clearly in Fig. 5. Here we again plot the spectra at the validation $\chi^{2}$ minimum (red) and at the temperature where $a=0.5$ (blue), as well as at a slightly lower/higher temperature, corresponding to one standard deviation below/above from the validation minimum (green/yellow). It is interesting to note that, at the lower sampling temperature (green), SAC is able to reproduce certain features of the exact spectrum better than at the temperatures corresponding to the validation minimum or $a=0.5$ (such as the position of the first peak and area where the two peaks intersect), while reproducing others not as well (such as the position of the second peak). This is consistent with the broadness of the validation $\chi^{2}$ minimum, as it is not necessarily clear which is the better choice of spectra. Furthermore, all of the spectra shown here are acceptable reproductions of the underlying artificial spectrum; the most prominent features of the spectrum are captured well in all four cases and the deviations plotted in Fig. 5 reflect only small shifts in the locations and amplitudes of the two Gaussian peaks.
@@ V.1 Two Equal Amplitude Gaussian Peaks
In the first case we consider, we used an artificial spectrum composed of two equal amplitude Gaussian peaks centered at $\omega=2$ and $\omega=4$. While this spectrum is relatively simple, the proximity of the two peaks poses similar difficulties to those encountered in Sec. IV.2; the low amplitude region in between the two peaks cannot be resolved fully at this error level, which causes a ripple effect of distortions throughout the spectrum. The two unconstrained sampling schemes, EA and VA, are able to overcome these challenges to different extents, which can be seen clearly when preforming tests using artificial spectra. However, when preforming analytic continuation using QMC-generated data, one must instead rely on cross validation to asses the abilities of these two parameterizations.
@@
The EA parameterization (panel (a)) produces a spectrum where both the shapes and heights of the two Gaussian peaks are very close to those in the exact spectrum, shown in black. On the other hand, the VA parameterization (panel (b)) produces peaks that are far too sharp with amplitudes that are unequal and too large. While both parameterizations are unable to correctly reproduce the region in between the two peaks, the VA parameterization preforms far worse. The excessively tall peaks produced by the variable amplitudes parameterization may be compensation for the missing weight in this region.
@@ VI Optimal Spectral Parameterization: Constrained Sampling
where $A_{0}+\sum_{i}A_{i}=1$ and the constraint $\omega_{i}>\omega_{0}$ for all $i$ is imposed during the sampling process. With this parameterization, the location of the peak, $\omega_{0}$, is sampled within the program, along with all of the other $\delta$-functions, but the weight of the macroscopic $\delta$-peak, $A_{0}$, is fixed and optimized using a simple scan to find a $\chi^{2}$ minimum, as detailed in Appendix A.1. This type of spectral function is appropriate for describing the spectrum of the $S=1/2$ AFM Heisenberg model in two-dimensions (2D) [28], where linear spin-wave theory predicts the presence of a macroscopic, “single-magnon” peak [29], while an incoherent continuum at higher energy can be attributed to either multimagnon excitations or partially deconfined spinon excitations (both of which cannot be captured by convention spin-wave theory [27]). The SAC parameterization in Eq. (19) has been used to produce results for the spectral function of the aforementioned square-lattice Heisenberg model that agreed very well with results from inelastic neutron-scattering experiments on a material considered the best physical realization of this model [27].
@@
In most cases, one tests both the unconstrained and constrained parameterizations when preforming analytic continuation on QMC-generated data. However, relying on the value of $\chi^{2}$ alone to compare the resulting spectra is not always reliable, since the SAC method can produce spectra with acceptable $\chi^{2}$ values even when the improper parameterization is being used (a consequence of the “ill-posed” problem). Our cross validation procedure can be used as an unbiased method not only to determine whether an unconstrained or constrained sampling scheme should be used, but also to determine which of the constrained sampling scheme produces the most statistically likely spectrum.
@@ VI.1 Heisenberg Chain
We first demonstrate the ability to discriminate between models using real QMC data, generated using the stochastic series expansion (SSE) method [33], for the $S=1/2$ AFM Heisenberg chain. We followed the procedure outlined in Sec. III, again with $K=20$ validation sets with an error level of $\sigma=10^{-5}$. The system size used here is $L=512$, and the QMC simulation was run using the inverse temperature $\beta=1024$, which is large enough to access ground state properties. We considered $S_{q}^{z}(\omega)$ at two momenta, $q=\pi/2$ and $43\pi/$ 4, as test cases. For both $q$ values, the optimal weights of the macroscopic $\delta$-peak were determined using the method detailed in Ref. 20 and shown here in Appendix A.1 ($A_{0}=0.40$ and $0.35$ for $q=\pi/2$ and $43\pi/$ 4, respectively). Due to the logarithmic correction to the square-root divergent edge form of the spectral function for $O=S_{q}^{z}$ [30, 31, 32], the optimal power $p$ for each $q$ value considered may not necessarily be equal to $-0.50$. As we show in Appendix A.2, for $q=3\pi/4$, the optimal exponent is $p=-0.60$.
@@
While it is difficult to glean any information about the relative performance of the three parameterizations from the sampling $\chi^{2}$, the validation $\chi^{2}$ paints a much different picture. At each temperature, the the edge parameterization has the lowest validation $\chi^{2}$, followed by the single-peak case, and then unconstrained sampling. This would suggest that the edge is indeed the most appropriate parameterization, in agreement with the BA solution for this model. This conclusion can also be made when comparing the values of the validation $\chi^{2}$ at fixed $\chi^{2}$. For $q=\pi/2$, the edge clearly preforms the best of the three parameterizations during the entire annealing run. For $q=3\pi/4$, the validation $\chi^{2}$ for the edge parametrization only dips below that of the other two at the end of the annealing run, but clearly reaches the lowest minimum value. We thus find that it is important take into consideration both the minimum value of each validation $\chi^{2}$ curve and the path the curves take when comparing the various parameterizations.
@@
The fact that the edge outperforms the unconstrained and single-peak parameterization strongly supports the viability of cross validation as a tool for model selection. Considering the fact that for a chain of length $512$, the exact $T=0$ spectrum likely only contains $\sim$100 delta functions with significant weight, deduced by extrapolating from the results for smaller system sizes [36, 37], it is remarkable that our method is able correctly identify the true form of the spectral function in the thermodynamic limit. It is quite possible that if smaller system size was considered instead, a spectrum composed of a dominate peak plus a smooth continuation would actually be the best representation and the single-peak parameterization would preform the best. But this is not the case here, implying that finite size effects are sufficiently mitigated.
@@ VII Discussion
A promising, and more decisive use of cross validation, is in model selection. The different entropy contents of the two unconstrained SAC parameterizations result in differing abilities to reproduce a given spectrum depending on its exact features. In some cases, it it advantageous to use the equal amplitudes parameterization, and in others the variable amplitudes parameterization preforms substantially better. In previous studies, this was only able to be shown retrospectively, when comparing the SAC results to an artificial spectrum used to generate synthetic data. Here, we showed that cross validation can indeed act as a tool for model selection when comparing the two different unconstrained SAC parameterizations. This will be extremely valuable in practice, as often times these two parameterizations produce considerably different results, which can be even more dramatic when other parameterizations are considered, such as the fixed grid method [20] or the orthogonal polynomial representation [41]. When there is no artificial spectrum to compare to, it is impossible to determine which of these spectra is best.
@@ Acknowledgements.
Appendix A Optimal SAC Parameters for the Heisenberg Chain
When implementing the single-peak and edge SAC parameterizations, a fixed parameter must be chosen using a simple optimization procedure. In the case of the single-peak parameterization, this parameter is the weight of the macroscopic $\delta$-peak, $A_{0}$, and in the case of the edge parameterization, this parameter is the power $p$ with which the edge asymptotically diverges with. Scans over $A_{0}$ and $p$ are preformed before cross validation to determine their optimal values.
In either case, the SAC annealing routine in run multiple times with different fixed values of $A_{0}$ or $p$, where typically the values are selected from a sensibly defined grid (not too dense or too sparse). After these scans are preformed, the $\chi^{2}$ values as a function of either $A_{0}$ or $p$ are compared at a fixed value of the sampling temperature and a minimum is located. At very low values of $\Theta$, where the sampling is dominated by noise in the QMC-generated data, the $\chi^{2}$ minimum will be very shallow, if present at all, so instead an elevated temperature is used. This temperature can be chosen based on the minimum $\chi^{2}$ reach in an unrestricted sampling SAC run on the same set of data, as detailed in Ref. 20, or simply set to a value high enough for a clear minimum to emerge (but low enough so that the $\chi^{2}$ value at the minimum is still acceptable). The exact location of the minimum depends on the sampling temperature to some extent, but in most cases this dependance is very minor. Furthermore, any value of $A_{0}$ or $p$ chosen around the $\chi^{2}$ minimum will produce nearly identical spectra.
It is crucial to have as highest possible data quality when determining the optimal values of $A_{0}$ and $p$. Therefore, we use the full set of QMC-generated data for the Heisenberg chain when preforming the parameter scans. In this case, the complete $G(\tau)$ data set has an error level $\sim~{}3~{}\times~{}10^{-6}$, nearly ten times smaller than the error level of the validation data sets.
A.1 $A_{0}$ Scan

### v3

@@ I Introduction
Dynamic response functions (spectral functions) provide a key link between the quantum Monte Carlo (QMC) simulations of many-body systems and their experimental counterparts. While it is intractable to measure these spectral function directly by QMC simulations, one can instead perform numerical analytic continuation of imaginary-time correlation functions to real frequency. However, this process is what is considered an “ill-posed” problem [1, 2, 3], meaning that there may be many viable solutions for the spectral function fitting a set of QMC-generated data. The Maximum Entropy method (MEM) [4, 5, 6, 7] has been the traditional tool used to tackle this problem and has produced many important results [8]. However, MEM has its limitations, being able only to reproduce broad spectral features, unless prior information about more complicated structure is provided to the algorithm.
@@
Cross validation has a natural extension to SAC. The imaginary-time correlation function data, calculated using QMC simulations, are typically stored as individual bins, which can be split up into sampling and validation sets. Here, the goodness-of-fit $\chi^{2}$ is used as the loss function to “train” and to “validate” the resulting spectral function. The concept of using cross validation within SAC was first explored by Efremkin et al. [24], whose work we draw on and significantly expand upon. Here, we perform an exhaustive study using both synthetic and real QMC-generated data.
In Sec. II we describe the SAC procedure generally, and in Sec. III we explain our cross validation scheme in detail. In Sec. IV.1 and  IV.2 we implement cross validation using the unconstrained SAC method and demonstrate how it can be used to determine the optimal sampling temperature, which we compare to the simple criterion proposed previously [20]. In Sec. V.1 and  V.2, we show how cross validation can be used in practice to determine which unconstrained SAC sampling parameterization produces the most statistically likely spectrum. We extend this model selection procedure to the class of constrained SAC parameterization in Sec. VI. In Sec. VI.1 we perform a test using QMC-generated data for the $S=1/2$ antiferromagnetic (AFM) Heisenberg model, where it is known that the spectral function for the operator $O=S_{q}^{z}$ contains a power-law divergent edge. In Sec. VI.2 we perform a preliminary test on the $S=1/2$ AFM Heisenberg model with long-range interactions [25, 26], a model where the exact features of the spectral function are unknown. We summarize our results and discuss the remaining open questions and future directions in Sec. VII.
@@ II SAC Procedure
$\bar{G}_{k}(\tau_{i})=\frac{1}{N_{k}}\sum_{b\in k}G^{b}(\tau_{i}),$ (6)
where the sum is over all bins $b$ in the $k$th validation data set. We define the “error level” for these sets of $G(\tau)$ data as the magnitude of the error bar on the averaged correlation function, $\bar{G}_{k}(\tau_{i})$, for the largest $\tau$ point included, when normalized so that $\bar{G}_{k}(0)=G^{b}(0)=1$. For larger $\beta$, a cutoff is typically introduced at the $\tau$ value where the relative error on $\bar{G}_{k}(\tau)$ is greater than 10%, which we would instead use to quantify the error level.
@@ III Cross Validation Procedure
When preforming cross validation, it is important to consider the error level of the individual sampling and validation data sets. The number $K$ must be chosen such that the assumptions made in deriving Eq. (16) are valid, but should be small enough to maintain a low enough error level. As displayed in the many examples investigated in Ref. 20, an error level of $\sigma\sim 10^{-5}$ is typically sufficient to accurately reproduce the underlying spectra. When the spectral function contains many fine features, such as collections of small peaks or sharp edges, a smaller error level may be necessary to accurately resolve said features. However, for the spectra we consider here, this is not the case, so we have used data with $\sigma\sim 10^{-5}$ in all of the examples shown in the following sections. For the tests using QMC-generated data, this restricted us to using $K=20$ validation data sets. For consistency, we have also used this number of validation sets for our tests using synthetic data, where we have direct control of both the error level and number of $G(\tau)$ bins.
In Appendix A, we test an alternative cross validation procedure that maintains as high a data quality as possible in both the sampling and validation sets, splitting the $G(\tau)$ data into just two mutually exclusive sets.
@@ IV.1 Single Gaussian Peak
in some cases with $A_{i}$ variable and sampled [11], as depicted in Fig. 1(b), but in this case we use fixed amplitudes $A_{i}=1/N_{\omega}$, Fig. 1(a). In this example, we use $N_{\omega}=5,000$ total $\delta$-functions, which is large enough to no longer detect the dependence of $S(\omega)$ on $N_{\omega}$.
@@ IV.2 Two Gaussian Peaks
We performed the same test as in Sec. IV.1, and present the results in Fig. 4. In this case, both the validation minimum and the temperature range corresponding to the criterion with $a$ ranging from 0.25 and 1 are much broader, but still align quite well with one another. The higher level of uncertainty is a reflection of the inability of the unconstrained sampling scheme to reproduce this spectrum nearly as well as in the previous case, which can be seen clearly in Fig. 5. Here we again plot the spectra at the validation $\chi^{2}$ minimum (red) and at the temperature where $a=0.5$ (blue), as well as at a slightly lower/higher temperature, corresponding to one standard deviation below/above from the validation minimum (green/yellow). It is interesting to note that, at the lower sampling temperature (green), SAC is able to reproduce certain features of the exact spectrum better than at the temperatures corresponding to the validation minimum or $a=0.5$ (such as the position of the first peak and area where the two peaks intersect), while reproducing others not as well (such as the position of the second peak). This is consistent with the broadness of the validation $\chi^{2}$ minimum, as it is not necessarily clear which is the better choice of spectra. Furthermore, all of the spectra shown here are acceptable reproductions of the underlying artificial spectrum; the most prominent features of the spectrum are captured well in all four cases and the deviations plotted in Fig. 5 reflect only small shifts in the locations and amplitudes of the two Gaussian peaks.
@@ V.1 Two Equal Amplitude Gaussian Peaks
In the first case we consider, we used an artificial spectrum composed of two equal amplitude Gaussian peaks centered at $\omega=2$ and $\omega=4$. While this spectrum is relatively simple, the proximity of the two peaks poses similar difficulties to those encountered in Sec. IV.2; the low amplitude region in between the two peaks cannot be resolved fully at this error level, which causes a ripple effect of distortions throughout the spectrum. The two unconstrained sampling schemes, EA and VA, are able to overcome these challenges to different extents, which can be seen clearly when performing tests using artificial spectra. However, when performing analytic continuation using QMC-generated data, one must instead rely on cross validation to asses the abilities of these two parameterizations.
@@
The EA parameterization (panel (a)) produces a spectrum where both the shapes and heights of the two Gaussian peaks are very close to those in the exact spectrum, shown in black. On the other hand, the VA parameterization (panel (b)) produces peaks that are far too sharp with amplitudes that are unequal and too large. While both parameterizations are unable to correctly reproduce the region in between the two peaks, the VA parameterization performs far worse. The excessively tall peaks produced by the variable amplitudes parameterization may be compensation for the missing weight in this region.
@@ V.2 Sharp Gaussian Peak with a Shoulder
At this point it should be pointed out that in the four unconstrained sampling tests we have presented, the fidelity of the SAC spectrum varies depending on the exact form of the underlying spectrum. In light of this, the cross validation $\chi^{2}$ value should not be taken as an absolute measure, but rather a relative statistic used to compare two (potentially imperfect) solutions. The ability of cross validation to identify which imperfect solution is more accurate, as gauged by direct comparison to the exact spectrum, is a testament to the robustness of this procedure.
@@ VI Optimal Spectral Parameterization: Constrained Sampling
where $A_{0}+\sum_{i}A_{i}=1$ and the constraint $\omega_{i}>\omega_{0}$ for all $i$ is imposed during the sampling process. With this parameterization, the location of the peak, $\omega_{0}$, is sampled within the program, along with all of the other $\delta$-functions, but the weight of the macroscopic $\delta$-peak, $A_{0}$, is fixed and optimized using a simple scan to find a $\chi^{2}$ minimum, as detailed in Appendix B.1. This type of spectral function is appropriate for describing the spectrum of the $S=1/2$ AFM Heisenberg model in two-dimensions (2D) [28], where linear spin-wave theory predicts the presence of a macroscopic, “single-magnon” peak [29], while an incoherent continuum at higher energy can be attributed to either multimagnon excitations or partially deconfined spinon excitations (both of which cannot be captured by convention spin-wave theory [27]). The SAC parameterization in Eq. (19) has been used to produce results for the spectral function of the aforementioned square-lattice Heisenberg model that agreed very well with results from inelastic neutron-scattering experiments on a material considered the best physical realization of this model [27].
@@
In most cases, one tests both the unconstrained and constrained parameterizations when performing analytic continuation on QMC-generated data. However, relying on the value of $\chi^{2}$ alone to compare the resulting spectra is not always reliable, since the SAC method can produce spectra with acceptable $\chi^{2}$ values even when the improper parameterization is being used (a consequence of the “ill-posed” problem). Our cross validation procedure can be used as an unbiased method not only to determine whether an unconstrained or constrained sampling scheme should be used, but also to determine which of the constrained sampling scheme produces the most statistically likely spectrum.
@@ VI.1 Heisenberg Chain
We first demonstrate the ability to discriminate between models using real QMC data, generated using the stochastic series expansion (SSE) method [33], for the $S=1/2$ AFM Heisenberg chain. We followed the procedure outlined in Sec. III, again with $K=20$ validation sets with an error level of $\sigma=10^{-5}$. The system size used here is $L=512$, and the QMC simulation was run using the inverse temperature $\beta=1024$, which is large enough to access ground state properties. We considered $S_{q}^{z}(\omega)$ at two momenta, $q=\pi/2$ and $43\pi/$ 4, as test cases. For both $q$ values, the optimal weights of the macroscopic $\delta$-peak were determined using the method detailed in Ref. 20 and shown here in Appendix B ($A_{0}=0.40$ and $0.35$ for $q=\pi/2$ and $43\pi/$ 4, respectively). Due to the logarithmic correction to the square-root divergent edge form of the spectral function for $O=S_{q}^{z}$ [30, 31, 32], the optimal power $p$ for each $q$ value considered may not necessarily be equal to $-0.50$. As we show in Appendix B.2, for $q=3\pi/4$, the optimal exponent is $p=-0.60$.
@@
While it is difficult to glean any information about the relative performance of the three parameterizations from the sampling $\chi^{2}$, the validation $\chi^{2}$ paints a much different picture. At each temperature, the the edge parameterization has the lowest validation $\chi^{2}$, followed by the single-peak case, and then unconstrained sampling. This would suggest that the edge is indeed the most appropriate parameterization, in agreement with the BA solution for this model. This conclusion can also be made when comparing the values of the validation $\chi^{2}$ at fixed $\chi^{2}$. For $q=\pi/2$, the edge clearly performs the best of the three parameterizations during the entire annealing run. For $q=3\pi/4$, the validation $\chi^{2}$ for the edge parametrization only dips below that of the other two at the end of the annealing run, but clearly reaches the lowest minimum value. We thus find that it is important take into consideration both the minimum value of each validation $\chi^{2}$ curve and the path the curves take when comparing the various parameterizations.
@@
The fact that the edge outperforms the unconstrained and single-peak parameterization strongly supports the viability of cross validation as a tool for model selection. Considering the fact that for a chain of length $512$, the exact $T=0$ spectrum likely only contains $\sim$100 delta functions with significant weight, deduced by extrapolating from the results for smaller system sizes [36, 37], it is remarkable that our method is able correctly identify the true form of the spectral function in the thermodynamic limit. It is quite possible that if smaller system size was considered instead, a spectrum composed of a dominate peak plus a smooth continuation would actually be the best representation and the single-peak parameterization would perform the best. But this is not the case here, implying that finite size effects are sufficiently mitigated.
To supplement these results, we have performed a second cross validation test in which the $G(\tau)$ bins are split into two mutually exclusive sets, instead of 21. Here, one set is used for sampling and the other is used for validation, but now we repeat the partitioning many times, preforming many runs of the cross validation procedure in a “bootstrap-style.” The validation $\chi^{2}$ value is then averaged over these repetitions and compared amongst the different SAC parameterizations. We performed this secondary test to ensure that the reduction of data quality, as a consequence of using only a fraction of the data for sampling, did not skew our results. The results agree with those shown in Fig. 10, which we present and discuss in Appendix A.
@@ VII Discussion
A promising, and more decisive use of cross validation, is in model selection. The different entropy contents of the two unconstrained SAC parameterizations result in differing abilities to reproduce a given spectrum depending on its exact features. In some cases, it it advantageous to use the equal amplitudes parameterization, and in others the variable amplitudes parameterization performs substantially better. In previous studies, this was only able to be shown retrospectively, when comparing the SAC results to an artificial spectrum used to generate synthetic data. Here, we showed that cross validation can indeed act as a tool for model selection when comparing the two different unconstrained SAC parameterizations. This will be extremely valuable in practice, as often times these two parameterizations produce considerably different results, which can be even more dramatic when other parameterizations are considered, such as the fixed grid method [20] or the orthogonal polynomial representation [41]. When there is no artificial spectrum to compare to, it is impossible to determine which of these spectra is best.
@@ Acknowledgements.
Appendix A Alternative Cross Validation Procedure
As briefly introduced in Sec. VI.1, we have performed an alternative cross validation procedure that strays slightly from the scheme introduced in the main text. In this approach, we maintain as high a data quality as possible in both the sampling and validation data sets by splitting the $G(\tau)$ bins into two mutually exclusive sets, instead of 21. To reduce fluctuations in the $\chi^{2}_{\mathrm{val}}$ statistic, we repeat this process many times, with new random partitions for the sampling and validation sets, and average the results (500 repetitions for the examples we present here).
Figure 14: Cross validation results for $S(q=\pi/2,\omega)$ and $S(q=3\pi/4,\omega)$ of the $S=1/2$ Heisenberg chain ($L=512$, $\beta=1024$) using the alternative cross validation procedure. Panels (a) and (b) show validation $\chi^{2}$ as a function of the sampling $\chi^{2}$ for $q=\pi/2$ and $q=3\pi/4$, respectively.
We tested this alternative procedure on the $S=1/2$ Heisenberg chain and present the results for for $S(q=\pi/2,\omega)$ and $S(q=3\pi/4,\omega)$ in Fig. 14. The same conclusions are drawn from both cross validation methods (cf. Fig. 10); the edge parameterization has the lowest validation $\chi^{2}$ when compared to the single-peak and unconstrained methods, at a fixed value of the $\chi^{2}$. The validation $\chi^{2}$ curves for the peak and edge parameterizations ($q=\pi/2$) and unconstrained sampling ($q=3\pi/4$) do not have well defined minima, which differs from the results in Fig. 10. This, along with the the other intricacies involved in using this approach, will be studied more closely in a future paper. Nevertheless, the results of this test are clearly inline with those presented in Sec. VI.1 and emphasize that the value of the validation $\chi^{2}$ throughout the entire anneal, as well as its minimum value, are important to consider when preforming cross validation.
This method offers several advantages, such as maintaining the highest possible data quality and the ability to preform many repetitions until the results have converged, but does not appear to produce results that conflict with our original approach. This is consistent with our experience using SAC. Increasing the error level by a factor of $\sqrt{21}\approx 4.5$ (which is the effect of using only 1/21st of the total number of $G(\tau)$ bins) will rarely be detrimental to the accurately of the results, assuming the data quality is high to begin with and the spectral function does not contain many fine features. In a future study, we will investigate this balance of data quality and number of validation data sets in more detail, but these preliminary results suggest that lowered data quality has not impacted our cross validation procedure.
Appendix B Optimal SAC Parameters for the Heisenberg Chain